Is T ! a concave, convex function? or neither? for σ E (Q2).
Answer #1

A function y = f(x) is convex (concave) if f''(x) = d2y/dx2 is positive (negative).

y = [c(1 - \sigma ) - 1] / (1 - \sigma )

y = [c(1 - \sigma ) / (1 - \sigma )] - [1 / (1 - \sigma )]

f'(c) = dy/dc = [1 / (1 - \sigma )] x [(1 - \sigma )] x c[(1 - \sigma ) - 1] = c (- \sigma )

f''(c) = d2y/dc2 = - \sigma x c(- \sigma - 1) = - \sigma / [c(\sigma + 1)]

Since 0 <= \sigma <= 2,

1 <= (\sigma + 1) <= 3

Therefore, f''(c) > 0, signifying that y = [c(1 - \sigma ) - 1] / (1 - \sigma ) is a convex function.